OpenStudy (kimberlybaker767):

Solve the system y = -x + 7 and y = -0.5(x - 3)2 + 8

OpenStudy (kimberlybaker767):

how do i do this

OpenStudy (radar):

There are several methods you may use. Substitution is a popular method and would be a good method because both equations equal y. Since y = y then the expressions to the right of the equal sign are also equal.

OpenStudy (radar):

This being the case we now can write: -x + 7 = -0.5(x - 3)2 + 8 In order to be sure I believe what is intended here is: \[-x + 7 = -0.5(x - 3)^{2} + 8\]Is the case?

OpenStudy (radar):

*Is this the case?

OpenStudy (kimberlybaker767):

yes

OpenStudy (kimberlybaker767):

Thank you for helping me.

OpenStudy (radar):

O.K. Firs clear the parentheses on the right expression: You should get a result of.... -x + 7 = -0.5(x^2 - 6x + 9) + 8 Now multiply thru by 2 getting -2x + 14 = -x^2 + 6x -9 + 16 Simplify the equation getting: x^2 -8x + 7= 0 Now you have a quadratic equation that can be easily factored and solve giving two values for x. Have you studied the solutions of quadratic equations yet?

OpenStudy (radar):

Factor this quadratic equation getting: (x - 7)(x - 1) = 0 Now solve getting: x - 7 = 0, thereby x = 7 and x - 1 = 0, thereby x = 1 two results, we are not thru, we got to solve for y. solve and verify by substituting these values in EACH of the original equations.

OpenStudy (radar):

Verification: y = -x + 7 for x = 1, y = -1 + 7 = 6 or (1,6) y - -x + 7 for x = 7, y = -7 + 7 = 0 or (7,0) Now for the critical test use the next original equation, if correct we will come out with same results.

OpenStudy (radar):

-0.5(x - 3)2 + 8\[y = -0.5(1 - 3)^{2} + 8 =-0,5(-2)^{2}+8=-2+8=6\] (1,6) same as first equation good so far, but has to work out with x = 7. \[y=-0.5(7 - 3)^{2} + 8 = -8 + 8 = 0\]Does indeed check out (7,0) x=7 y=0. That is the solutions of that system X=1, Y=6 and X=7, Y=0

OpenStudy (radar):

Did you follow, if not say where you went off the rail.

OpenStudy (radar):

I will wait for your response before logging off.

OpenStudy (kimberlybaker767):

i understand it all thank you again for helping me

OpenStudy (radar):

You're welcome and good luck with your studies.